# Video-RAG: Visually-aligned Retrieval-Augmented Long Video Comprehension

### v3

@@ 4.5 Qualitative Evaluation
We present qualitative results in the case of Video-MME [6] in Figure 5 and Figure 6. As illustrated, augmenting LLaVA-Video with external tools to process and retrieve auxiliary texts from videos significantly enhances its ability to reduce visual hallucinations, thereby enabling more accurate responses to user queries. Grad-CAM [32] and t-SNE [37] visualization results also show that applying Video-RAG helps the LVLM’s cross-modality alignment.
@@ 5 Conclusion
In this paper, we present Video-RAG for effective long video understanding by integrating retrieved auxiliary texts with LVLMs, achieving proprietary-level performance with 72B open-source LVLM. Unlike traditional long-context and GPT-based Agent methods that may have limited performance gain and are resource-intensive, Video-RAG offers a resource-efficient, plug-and-play solution for any open-source LVLMs that only leverage open-source external tools to extract the visually-aligned auxiliary texts from pure video data for input. In the future, we will explore how to more efficiently integrate auxiliary texts and provide an adaptive frame selection strategy for LVLMs.
Supplementary Material
6 Decouple Query
@@
7 Sub-set of Video-MME
As outlined in the implementation details, we randomly sampled a subset of the Video-MME [6] dataset to evaluate a computationally resource-intensive, agent-based method with long-context LVLMs. Specifically, we selected 10% of the full dataset, comprising 30 short, 30 medium-length, and 30 long videos. Each video contains three multiple-choice questions. Importantly, we ensured that the performance ranking of the methods on the subset mirrored that of the full dataset. As shown in Tables 6 and 7, we evaluated four distinct 7B models Chat-Univi-v1.5 [10], LLaVA-NeXT-Video [48], LongVA [47], and Long-LLaVA [43] using a frame sampling rate of 16 for both the subset and the full set. Our results indicate that the performance rankings remained consistent across both evaluations.
Method Short Medium Long Overall
Chat-Univi-v1.5 [10] 50.0 33.3 17.8 33.7
LLaVA-NeXT-Video [48] 54.4 33.3 23.3 37.0
LongVA [47] 56.7 50.0 38.9 48.5
Long-LLaVA [43] 58.9 52.2 40.0 50.4
@@ 7 Sub-set of Video-MME
Chat-Univi-v1.5 [10] 45.7 39.0 35.7 40.1
LLaVA-NeXT-Video [48] 51.1 41.8 36.8 43.2
LongVA [47] 60.8 45.2 41.4 49.1
Long-LLaVA [43] 59.3 49.3 44.4 51.0
@@
8 Results on Video-MME Sub-Set
We examine Video-RAG against two representative methods in terms of inference time, GPU resource requirements, and overall performance. Given that GPT-based Agent methods are resource-intensive, we randomly sampled a sub-set of the Video-MME [6] for evaluation, as described in Section 7. As demonstrated in Figure 7, VideoAgent [4], a typically GPT-based Agent method, requires significant time to process video and deliver suboptimal performance. Meanwhile, LongVA [47], a representative long-context LVLM, shows limited improvement from increasing the frame rate and even experiences performance degradation. Integrating our Video-RAG into the 16-frame LongVA results in substantial performance improvements while reducing GPU resource consumption. Specifically, with only increasing 8GB GPU memory compared to the base (16-frames LongVA), we achieve 11.5% overall performance improvement, while outperforming another long-context LVLM Long-LLaVA-7B [43] in 16-frames setting by 9.6% with less GPU memory requirements and compatible total inference time. These results demonstrated that our Video-RAG is lightweight with lower computing overhead than the other typical methods.
Figure 7: The comparison of our Video-RAG with two common approaches. The size of the bubbles represents the total time consumed for completing inference on the Video-MME [6] sub-set.
9 Details of Similarity Score Calculation
In the process of using the RAG system to retrieve auxiliary texts extracted from videos, we define a similarity threshold $t$ to ensure the selection of relevant texts. Specifically, we employ FAISS-based [11] similarity to select OCR and ASR texts, while CLIP [29] similarity is used for keyframe selection. In our implementation, the similarity threshold $t$ is set to 0.3. As for OCR and ASR selection, For any given list of the retrieve request $\bm{\mathrm{R}}$ and auxiliary texts $\bm{\mathrm{A}}$, the Contriever [8] framework maps the text to a text embedding as:
@@ 9 Details of Similarity Score Calculation
As for object detection, we use CLIP to select the video keyframe. During this process, we first filter the object detection request $\bm{\mathrm{R}}_{det}$ to ensure they correspond to CLIP-sensitive physical entities, avoiding the inclusion of abstract concepts. Specifically, if it is a single word, direct part-of-speech filtering is applied; if it is a compound word, certain rules are followed to check for compliance, such as whether it is an adjective plus a noun, or a noun plus a noun. We use the Spacy library to achieve this. After this, we put the text “A picture of” before each object detection request.
@@
[…]
@@
10 More Ablation Studies
Effect of different components of Video-RAG. We evaluate the performance across sub-tasks within Video-MME [6], as shown in Figure 8. The results reveal that object detection auxiliary texts significantly enhance spatial perception and object counting, while OCR auxiliary texts specifically improve performance on text recognition tasks. Additionally, ASR auxiliary texts contribute to a general improvement in inference tasks, underscoring the critical role of audio transcription in video understanding. Given that audio transcription is considerably more time-consuming than character recognition or object detection, these texts should be selected based on the requirements of the application.
Figure 8: Performance on 12 sub-tasks in Video-MME [6] benchmark after applying different components in Long-LLaVA.
Besides studying the inference of different components of Video-RAG in the Video-MME [6] benchmark, we also experiment with a different type of video benchmark VNBench [51] with Long-LLaVA-7B [43]. VNBench is a synthetic benchmark designed to evaluate models’ long-context abilities, covering tasks such as retrieval, ordering, and counting. VNBench randomly inserts stickers or text into the video that has nothing to do with the original content of the video, thus typically challenging the model’s needle-in-the-haystack capability. As shown in Table 8, we find that applying DET and OCR as auxiliary texts can significantly improve the performance in retrieval, ordering, and counting tasks. However, the ASR component will decline the performance due to the subtitles are not ancillary to this particular task. These results demonstrated that our proposed distinct types of auxiliary texts can be selected according to the application needs to meet the requirements better.
@@ 10 More Ablation Studies
Table 8: Results on combinations of different auxiliary texts in VNBench [51] with 1-try setting when applying 7B Long-LLaVA [43] as LVLM under the 32-frames setting. Ret, Ord, and Cnt represent retrieval, ordering, and counting tasks, respectively.
11 More Qualitative Results
In this section, we show more results of LLaVA-Vdieo-7B when applying Video-RAG in different examples in Figure 10. The figure highlights several representative cases involving detailed video comprehension from Video-MME [6]. As illustrated, augmenting LLaVA-Video with external tools to process and retrieve auxiliary texts from videos significantly enhances its ability to reduce visual hallucinations, thereby enabling more accurate and confident responses to user queries.
Figure 9: Decouple prompt of the multiple-choice question for LVLMs.

### v4

@@ 4.5 Qualitative Evaluation
We present qualitative results in the case of Video-MME [videomme] in Figure 4 and Figure 5. As illustrated, augmenting LLaVA-Video with external tools to process and retrieve auxiliary texts from videos significantly enhances its ability to reduce visual hallucinations, thereby enabling more accurate responses to user queries. Grad-CAM [gradcam] and t-SNE [tsne] visualization results also show that applying Video-RAG helps the LVLM’s cross-modality alignment.
@@ 5 Conclusion
In this paper, we present Video-RAG for effective long video understanding through integrating retrieved auxiliary texts with LVLMs, achieving proprietary-level performance with 72B open-source LVLM. Unlike traditional methods that are resource-intensive with limited gains, Video-RAG offers a resource-efficient, plug-and-play solution leveraging only open-source tools to extract visually-aligned auxiliary texts from video data. However, Video-RAG may be limited by the visual tools we choose and their performance, which lacks adaptation. In the future, we will explore how to more efficiently integrate auxiliary texts and provide an adaptive frame selection strategy for LVLMs.
6 Acknowledge
This work was supported by the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. U21A20472, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001), and the Fundamental Research Funds for the Central Universities.
Supplemental Material
Appendix A Decouple Query
@@
Appendix B Sub-set of Video-MME
As outlined in the implementation details, we randomly sampled a subset of the Video-MME [videomme] dataset to evaluate a computationally resource-intensive, agent-based method with long-context LVLMs. Specifically, we selected 10% of the full dataset, comprising 30 short, 30 medium-length, and 30 long videos. Each video contains three multiple-choice questions. Importantly, we ensured that the performance ranking of the methods on the subset mirrored that of the full dataset. As shown in Tables 6 and 7, we evaluated four distinct 7B models Chat-Univi-v1.5 [chatunivi], LLaVA-NeXT-Video [llavanextvideo], LongVA [longva], and Long-LLaVA [longllava] using a frame sampling rate of 16 for both the subset and the full set. Our results indicate that the performance rankings remained consistent across both evaluations.
@@ Appendix B Sub-set of Video-MME
Chat-Univi-v1.5 [chatunivi] 50.0 33.3 17.8 33.7
LLaVA-NeXT-Video [llavanextvideo] 54.4 33.3 23.3 37.0
LongVA [longva] 56.7 50.0 38.9 48.5
Long-LLaVA [longllava] 58.9 52.2 40.0 50.4
@@
Method Short Medium Long Overall
Chat-Univi-v1.5 [chatunivi] 45.7 39.0 35.7 40.1
LLaVA-NeXT-Video [llavanextvideo] 51.1 41.8 36.8 43.2
LongVA [longva] 60.8 45.2 41.4 49.1
Long-LLaVA [longllava] 59.3 49.3 44.4 51.0
Appendix C Results on Video-MME Sub-Set
We examine Video-RAG against two representative methods in terms of inference time, GPU resource requirements, and overall performance. Given that GPT-based Agent methods are resource-intensive, we randomly sampled a sub-set of the Video-MME [videomme] for evaluation, as described in Section B. As demonstrated in Figure 6, VideoAgent [videoagent], a typically GPT-based Agent method, requires significant time to process video and deliver suboptimal performance. Meanwhile, LongVA [longva], a representative long-context LVLM, shows limited improvement from increasing the frame rate and even experiences performance degradation. Integrating our Video-RAG into the 16-frame LongVA results in substantial performance improvements while reducing GPU resource consumption. Specifically, with only increasing 8GB GPU memory compared to the base (16-frames LongVA), we achieve 11.5% overall performance improvement, while outperforming another long-context LVLM Long-LLaVA-7B [longllava] in 16-frames setting by 9.6% with less GPU memory requirements and compatible total inference time. These results demonstrated that our Video-RAG is lightweight with lower computing overhead than the other typical methods. Moreover, we provide detailed time consuming to construct three types of databases (which can be built in parallel) and inference per query, as shown in Table 8.
Table 8: Overall performance, databases construct and average inference time (include building databases) per query (#Time) in Video-MME-mini.
Model ASR OCR DET Total Time w/o subs w/ Video-RAG
#Time Overall #Time Overall
VideoAgent - - - - 14min 47.7 - -
LongVA-16fs 21min 2min 3min max(21, 2, 3)=21min 1s 48.5 1s + 5s 60.0
LongVA-128fs 21min 16min 16min max(21, 16, 16)=21min 8s 54.1 8s + 5s 63.3
LongVA-384fs 42min 48min 24min max(42, 48, 24)=48min 20s 53.7 20s + 11s 63.6
Figure 6: The comparison of our Video-RAG with two common approaches. The size of the bubbles represents the total time consumed for completing inference on the Video-MME [videomme] sub-set.
Appendix D Details of Similarity Score Calculation
In the process of using the RAG system to retrieve auxiliary texts extracted from videos, we define a similarity threshold $t$ to ensure the selection of relevant texts. Specifically, we employ FAISS-based [faiss] similarity to select OCR and ASR texts, while CLIP [clip] similarity is used for keyframe selection. In our implementation, the similarity threshold $t$ is set to 0.3. As for OCR and ASR selection, For any given list of the retrieve request $\bm{\mathrm{R}}$ and auxiliary texts $\bm{\mathrm{A}}$, the Contriever [contriever] framework maps the text to a text embedding as:
@@ Appendix D Details of Similarity Score Calculation
As for object detection, we use CLIP to select the video keyframe. During this process, we first filter the object detection request $\bm{\mathrm{R}}_{det}$ to ensure they correspond to CLIP-sensitive physical entities, avoiding the inclusion of abstract concepts. Specifically, if it is a single word, direct part-of-speech filtering is applied; if it is a compound word, certain rules are followed to check for compliance, such as whether it is an adjective plus a noun, or a noun plus a noun. We use the Spacy library to achieve this. After this, we put the text “A picture of" before each object detection request.
@@
$\text{Keyframe:}\quad\bm{\mathrm{F}}_{j}\quad\text{if}\quad S_{j}^{\text{norm}}>t$
@@
Appendix E More Ablation Studies
Effect of different components of Video-RAG. We evaluate the performance across sub-tasks within Video-MME [videomme], as shown in Figure 7. The results reveal that object detection auxiliary texts significantly enhance spatial perception and object counting, while OCR auxiliary texts specifically improve performance on text recognition tasks. Additionally, ASR auxiliary texts contribute to a general improvement in inference tasks, underscoring the critical role of audio transcription in video understanding. Given that audio transcription is considerably more time-consuming than character recognition or object detection, these texts should be selected based on the requirements of the application.
Figure 7: Performance on 12 sub-tasks in Video-MME [videomme] benchmark after applying different components in Long-LLaVA.
Besides studying the inference of different components of Video-RAG in the Video-MME [videomme] benchmark, we also experiment with a different type of video benchmark. We first evaluate LLaVA-Video in MLVU [mlvu] and LongVideoBench [lvb] in both 7B and 72B scale with the 64-frame setting, results are shown in Table 9. As demonstrated, when all components are combined, we get optimal performance in both datasets, including 7B and 72B scales. Specifically, the performance in MLVU [mlvu] even declined when the RAG system was not implemented.
Table 9: Ablation study in MLVU and LongVideoBench.
RAG DET OCR ASR 7B 72B
MLVU LVB MLVU LVB
70.8 56.6 73.1 61.9
71.0 56.5 73.4 63.2
71.3 56.8 73.5 63.4
72.4 58.7 73.8 65.4
70.3 58.3 72.9 64.0
Figure 8: Decouple prompt of the multiple-choice question for LVLMs.
Then, to better point out the role of DET and OCR, we evaluate Video-RAG in VNBench [vnbench] with Long-LLaVA-7B [longllava]. VNBench is a synthetic benchmark designed to evaluate models’ long-context abilities, covering tasks such as retrieval, ordering, and counting. VNBench randomly inserts stickers or text into the video that has nothing to do with the original content of the video, thus typically challenging the model’s needle-in-the-haystack capability. As shown in Table 10, we find that applying DET and OCR as auxiliary texts can significantly improve the performance in retrieval, ordering, and counting tasks. However, the ASR component will decline the performance due to the subtitles are not ancillary to this particular task. These results demonstrated that our proposed distinct types of auxiliary texts can be selected according to the application needs to meet the requirements better.
Table 10: Results on combinations of different auxiliary texts in VNBench [vnbench] with 1-try setting when applying 7B Long-LLaVA [longllava] as LVLM under the 32-frames setting. Ret, Ord, and Cnt represent retrieval, ordering, and counting tasks, respectively.
@@ Appendix E More Ablation Studies
Appendix F More Qualitative Results
@@
In this section, we show more results of LLaVA-Vdieo-7B when applying Video-RAG in different examples in Figure 9. The figure highlights several representative cases involving detailed video comprehension from Video-MME [videomme]. As illustrated, augmenting LLaVA-Video with external tools to process and retrieve auxiliary texts from videos significantly enhances its ability to reduce visual hallucinations, thereby enabling more accurate and confident responses to user queries.